OpenStudy (volpina):

Find the absolute maximum value for the function f(x) = x^2 - 3, on the interval [-4, 0) U (0, 3]. (2 points) 3 4 13 No maximum exists because f is not continuous Idk how to do this, plz include steps on how to get the answer, thank you! :)

OpenStudy (3mar):

"the absolute maximum value for a function" could be calculated from its first derivative by equating the first derivative to zero and find the zeros of this derived function!

OpenStudy (volpina):

So the first derivative would be -32t and if it equals 0, t would be equal 0 so the answer would be D?

OpenStudy (3mar):

"would be -32t" what \(t\) ? \[\Huge f(x) = x^2 - 3\] that is our equation................................

OpenStudy (volpina):

>.> i'm confused o-o

OpenStudy (3mar):

Why? The given function is \[\Huge f(x) = x^2 - 3\] and its first derivative is \[\Huge f'(x) = \frac{ d }{ dx }(x^2 - 3)=2x\]

OpenStudy (volpina):

sorry i'm confusing questions >.< but how do you find a max on a line?

OpenStudy (volpina):

oh wait... nvm heh it doesn't exist cuz it's a line

OpenStudy (3mar):

Let me simplify it to you.. you calculated \(f'(x)\) and it is \(2x\),,,what did you get after equating it to zero?

OpenStudy (volpina):

that would be 2x=0 so x=0

OpenStudy (3mar):

You would get that \[2x=0\\x=0\] this value of x=0 means that the slope of the function is undefined or zero.. "Critical points of f are when f'(x)=0 or when f'(x) is undefined" i.e x=0 is the critical point of f(x).

OpenStudy (volpina):

So the answer is D?

OpenStudy (3mar):

Step by step.... After getting that critical point \(x=0\), search if it is in the specified domain or not. Is in is or out of it?

OpenStudy (3mar):

@volpina \(x=0\) lies in the given domain or it is not included?

OpenStudy (volpina):

It's included because both ranges include 0

OpenStudy (3mar):

[-4, 0) U (0, 3] does that mean \(x=0\) included or not?

OpenStudy (volpina):

Wait, it's not because zero doesn't have a square bracket

OpenStudy (3mar):

Yes, That is what I want you to take care about.